# Preference-based Reinforcement Learning beyond Pairwise Comparisons: Benefits of Multiple Options

### v1

@@ 5.1 Suboptimality Gap of M-AUPO
Under the same setting as Theorem 1, let $\kappa:=\frac{e^{-4B}}{4}$. Then, with probability at least $1-\delta$, M-AUPO (Algorithm 3) achieves the following suboptimality gap:
@@
Discussion of Theorem 2. For sufficiently large $T$, the suboptimality gap in Theorem 2 matches the leading-order term of Theorem 1, while its second (non-leading) term is tighter by a factor of $\mathcal{O}(K^{2})$. However, the per-round computational cost of the RB parameter update is $K$ times higher than that of the PL parameter update (see Remark 1). Despite this, the result is particularly notable as it offers a rigorous theoretical explanation for the empirical success of RLHF in LLMs (e.g., Ouyang et al. [59]), where ranking feedback is decomposed into pairwise comparisons for parameter estimation.
@@ 5.2 Lower Bound
In this subsection, we derive a lower bound for our setting: PbRL with linear rewards under ranking feedback generated by a Plackett-Luce (PL) model. The proof is deferred to Appendix G.
@@
Suppose $T\geq d^{2}/(8K^{2})$. Define the feature space as $\Phi:=\mathcal{S}^{d-1}$, the unit sphere in $\mathbb{R}^{d}$, and let the parameter space be $\Theta=\{-\mu,\mu\}^{d}$, where $\mu=\sqrt{d/(8K^{2}T)}$. Then, for any policy $\widehat{\pi}_{T}\in\triangle_{\Phi}$ returned after collecting $T$ samples (using any sampling policy), the expected suboptimality gap is lower bounded as:
@@
Discussion of Theorem 3. Theorem 3 provides theoretical support for our upper bounds, particularly with respect to the dependency on $K$. Compared to the upper bounds in Theorems 1 and 2, the remaining gap is only a factor of $\frac{1}{\sqrt{K}}$. Closing this gap remains an open problem for future work. To the best of our knowledge, this is the first lower bound on the suboptimality gap that incorporates PL ranking feedback in PbRL and formally shows that the suboptimality gap can diminish as $K$ grows, highlighting the advantage of utilizing ranking feedback over simple pairwise comparisons.
@@ 6 Numerical Experiments
We conduct two sets of experiments to empirically validate our theoretical findings: (i) one using synthetic data (Subsection 6.1), and (ii) another using two real-world datasets (Subsection 6.2). We compare our proposed algorithm, M-AUPO, against three baselines: (i) DopeWolfe [79], which selects $K$ actions in a non-contextual setting; (ii) Uniform, which uniformly samples assortments of size $K$ at random; and (iii) Best&Ref constructs an action pair ($|S_{t}|=2$) by combining the action that maximizes the current reward estimate with another sampled from a reference policy (e.g., uniform random or SFT), following the setup in Online GSHF [89] and XPO [88]. In our experiments, the reference policy for Best&Ref is set to the uniform random policy.
@@ 6.1 Synthetic Data
The first two plots in Figure 1 show the suboptimality gap of M-AUPO under both the PL loss (3) and RB loss (4) as the maximum assortment size $K$ varies. The results clearly show that performance improves as $K$ increases, supporting our theoretical findings. In the third plot of Figure 1, we compare the performance of M-AUPO with other three baselines under the PL loss with $K=5$ at the final round, demonstrating that our algorithm outperforms other baselines significantly. While DopeWolfe also considers the selection of $K$ actions from $N$ actions, it treats each context $x$ independently and is specifically designed for the context-free setting (i.e., a singleton context). As a result, DopeWolfe cannot leverage information sharing across varying contexts and performs poorly in our setting. Furthermore, M-AUPO outperforms naive assortment selection strategies such as Uniform and Best&Ref, as it explicitly chooses assortments that maximize the expected uncertainty, thereby achieving more efficient exploration. See Appendix I.1 for additional experimental details and results.
@@ 6.2 Real-World Dataset
We use the gemma-2b [^5] https://huggingface.co/google/gemma-2b-it [78] LLM to construct the feature $\phi(x,a)$. Specifically, $\phi(x,a)$ is obtained by extracting the embedding of the concatenated prompt and response from the last hidden layer of the LLM, with size $d=2048$. Additionally, we use the Mistral-7B [33] reward model [^6] https://huggingface.co/Ray2333/reward-model-Mistral-7B-instruct-Unified-Feedback as the true reward model $r_{\bm{\theta}^{\star}}$ to generate ranking feedback and compute the suboptimality gap accordingly. We measure the suboptimality gap every 2,500 rounds and report the average over 10 independent runs, along with the standard error. In these experiments, we present only the results under the PL loss, as the performance difference between the PL and RB losses is negligible, as shown in Figure 1.
The first two plots in Figure 2 show the suboptimality gap of M-AUPO under the PL loss on two real-world datasets as the maximum assortment size $K$ varies. Consistent with our theoretical findings, the performance improves as $K$ increases. In the third plot of Figure 2, we compare the performance of M-AUPO with other baselines under the PL loss with $K=3$ at the final round, showing that M-AUPO outperforms baselines by a large margin, consistent with the results from the synthetic data experiment. See Appendix I.2 for additional experimental details and results.
Figure 1: Synthetic data experiment: suboptimality gap of M-AUPO under varying $K$, evaluated with PL loss (left) and RB loss (middle), along with comparison against DopeWolfe (right).
Figure 2: Real-world dataset experiment: suboptimality gap of M-AUPO under varying $K$ on the TREC-DL dataset (left) and the NECTAR dataset (middle), along with comparison against DopeWolfe (right). The results are rescaled to align the performances between the two datasets.
@@ 7 Conclusion
To the best of our knowledge, this work presents the first theoretical result in online PbRL showing that the suboptimality gap decreases as more options are revealed to the labeler for ranking feedback. By demonstrating its statistical efficiency, our results provide a solid theoretical foundation for moving beyond the prevalent reliance on pairwise comparisons. We hope this finding will encourage future research to explore richer feedback formats beyond pairwise comparisons.
Moreover, our analysis eliminates the $\mathcal{O}(e^{B})$ dependency in the leading term without introducing any additional algorithm. This result implies that all existing PbRL and dueling bandit algorithms can likewise avoid this harmful dependency without modification—indicating that the limitation lies in their analyses rather than in the optimality of the algorithms themselves. The key takeaway is that in PbRL, the $\mathcal{O}(e^{B})$ dependency is theoretically avoidable and thus no longer poses a limitation.
We believe that these two implications are both conceptually significant and provide meaningful contributions toward a deeper theoretical understanding of PbRL.
@@ Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2022-NR071853, RS-2023-00222663, and RS-2025-25420849), by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2025-02263754), and by AI-Bio Research Grant through Seoul National University.
@@ Appendix A Further Related Work
Logistic and MNL bandits. Our work is also closely related to logistic bandits and multinomial logit (MNL) bandits. The logistic bandit problem [23, 26, 2, 27, 44, 45] is a special case of the MNL bandit model in which the agent offers only a single item (i.e., $K=1$) at each round and receives binary feedback indicating whether the item was selected (1) or not (0). Faury et al. [26] examined how the regret in logistic bandits depends on the non-linearity parameter $\kappa$ of the logistic link function and proposed the first algorithm whose regret bound eliminates explicit dependence on $1/\kappa=\mathcal{O}(e^{B})$. Abeille et al. [2] further improved the theoretical dependency on $1/\kappa$ and established a matching, problem-dependent lower bound. Building on this, Faury et al. [27] developed a computationally efficient algorithm whose regret still matches the lower bound established by Abeille et al. [2].
@@
Multinomial logit (MNL) bandits tackle a more sophisticated problem than logistic bandits. Instead of offering a single item and observing binary feedback, the learner chooses a subset of items—underscoring the combinatorial nature of the task—and receives non-uniform rewards driven by an MNL choice model [5, 4, 58, 15, 55, 56, 61, 3, 41, 43]. A recent breakthrough by Lee and Oh [41] closed a long-standing gap by providing a computationally efficient algorithm that attains the minimax-optimal regret for this setting. Building on this result, Lee and Oh [43] further reduced the regret bound by a factor polynomial in $B$ and logarithmic in $K$, and established the first variance-dependent regret bounds for MNL bandits.
@@
Our work extends the online confidence bound analysis of Lee and Oh [43] to the Plackett–Luce (PL) model. This extension is natural because the PL probability distribution decomposes into a sequence of MNL probabilities over successive choices. Crucially, we leverage their key insight—that the MNL loss exhibits an $\ell_{\infty}$-self-concordant property—to eliminate the harmful $\mathcal{O}(e^{B})$ dependence. This is one of the main contributions of our work (see Lemma D.2).
@@
RL with MNL models. Recent work has extended the Multinomial Logit (MNL) framework beyond bandit formulations to reinforcement learning. Lee and Oh [42] introduced combinatorial RL with preference feedback, a framework in which an agent learns to select subsets of items so as to maximize long-term cumulative rewards.
@@
Another line of research incorporates MNL models directly into the transition dynamics. Hwang and Oh [29] proposed MNL-MDPs, a class of Markov decision processes whose transition probabilities follow an MNL parameterization. Building upon this formulation, Cho et al. [17] improved the regret bounds by improving the exponential dependence on $B$, and Park et al. [60] extended the analysis to the infinite-horizon setting.
Procedure A.1 Greedy Selection of Reference Action and Assortment
1:Input: $x_{t},H_{t}^{-1},\mathcal{A},K$
2:Initialize: $(\bar{a}_{t}^{\star},S_{t}^{\star},\text{max\_avg})\leftarrow(\text{None},\text{None},-\infty)$
3:for all $\bar{a}\in\mathcal{A}$ do
4:  Initialize $S\leftarrow\{\bar{a}\}$, $\text{prev\_avg}\leftarrow 0$
5:  while $|S|<K$ do
6:   Find
$a^{\star}\leftarrow\mathop{\mathrm{argmax}}_{a\in\mathcal{A}\setminus S}\left\|\phi(x_{t},a)-\phi(x_{t},\bar{a})\right\|_{H_{t}^{-1}}$
7:   Tentatively update $S^{\prime}\leftarrow S\cup\{a^{\star}\}$
8:   Compute
$\text{cur\_avg}\leftarrow\frac{1}{|S^{\prime}|}\sum_{a\in S^{\prime}}\left\|\phi(x_{t},a)-\phi(x_{t},\bar{a})\right\|_{H_{t}^{-1}}$
9:   if $\text{cur\_avg}<\text{prev\_avg}$ then
10:     break
11:   else
12:     $S\leftarrow S^{\prime}$
13:     $\text{prev\_avg}\leftarrow\text{cur\_avg}$
14:   end if
15:  end while
16:  if $\text{prev\_avg}>\text{max\_avg}$ then
17:   $(\bar{a}_{t}^{\star},S_{t}^{\star},\text{max\_avg})\leftarrow(\bar{a},S,\text{prev\_avg})$
18:  end if
19:end for
20:return $(\bar{a}_{t}^{\star},S_{t}^{\star})$
Appendix B Efficient Assortment Selection
In this section, we describe how the assortment selection rule in Equation (9) can be solved efficiently.
Given $x_{t}$, the reference action–assortment pair $(\bar{a}_{t},S_{t})$ is selected by evaluating each candidate reference action $\bar{a}\in\mathcal{A}$. For each $\bar{a}$, we construct an assortment $S$ beginning with the singleton $\bar{a}$}, and iteratively add actions $a\in\mathcal{A}\setminus\{\bar{a}\}$ in decreasing order of their uncertainty relative to $\bar{a}$, measured by
$\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}\right)\right\|_{H_{t}^{-1}}.$
Let $a_{tk}(\bar{a})$ denote the action with the $k$-th highest uncertainty with respect to $\bar{a}$ at round $t$. For example, $a_{t1}(\bar{a})=\mathop{\mathrm{argmax}}_{a\in\mathcal{A}\setminus\{\bar{a}\}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}\right)\right\|_{H_{t}^{-1}}$. We add actions greedily to the set $S$, as long as the average uncertainty continues to increase:
$\frac{1}{|S|}\sum_{a\in S}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}\right)\right\|_{H_{t}^{-1}},\quad\text{where }\,\bar{a}\in S.$
Among all candidates $\bar{a}\in\mathcal{A}$, we select the pair $(\bar{a}_{t},S_{t})$ that achieves the highest average uncertainty. The pseudocode is given in Procedure A.1.
For each candidate reference action, the algorithm incrementally constructs a subset of actions by greedily adding those with the highest uncertainty relative to the reference—stopping once the average uncertainty no longer increases. This greedy strategy guarantees that, for each reference, the selected subset maximizes the average uncertainty. By applying this procedure across all possible reference actions and selecting the pair that achieves the highest score, the algorithm obtains the global optimum over all reference–assortment combinations.
As for the computational cost, each greedy addition step involves searching over $\mathcal{O}(N)$ candidate actions, resulting in a total of $\tilde{\mathcal{O}}(NK)$ operations per each reference action $\bar{a}$. Repeating this process for all $N$ candidate references yields a total cost of $\tilde{\mathcal{O}}(N^{2}K)$.
Appendix C Notation
@@
Let $T$ denote the total number of rounds, with $t\in[T]$ representing the current round. We use $N$ for the total number of items, $K$ for the maximum assortment size, $d$ for the feature vector dimension, and $B$ as an upper bound on the norm of the unknown parameter. For notational convenience, we provide Table C.1.
@@
${\mathcal{T}}^{w}$ $\left\{t\in[T]:\max_{a\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|_{H_{t}^{-1}}\geq\frac{1}{3\sqrt{2}\beta_{T+1}(\delta)}\right\}$, warm-up rounds
$\Lambda_{t}$ $\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{a\in S_{s}}\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}$
${\mathcal{T}}_{0}$ $\left\{t\in[T]:\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{\Lambda_{t}^{-1}}\geq 1.\right\}$, large EP rounds
@@
Appendix D Proof of Theorem 1

### v2

@@ 5.1 Suboptimality Gap of M-AUPO
Under the same setting as Theorem 1, if the parameter is updated and the assortment is selected according to the RB loss setting, then for any $T\geq e^{K^{2}/d}$, with probability at least $1-\delta$, M-AUPO (Algorithm 3) satisfies
@@
Discussion of Theorem 2. For sufficiently large $T$, the suboptimality gap in Theorem 2 matches the leading-order term of Theorem 1, while its second (non-leading) term is tighter by a factor of $\mathcal{O}(K^{4})$. In addition, the assortment selection cost for RB (12) is lower, since it avoids the expensive expectation over rankings required in the PL loss case (Equation (10)). This leads to a useful insight: under ranking feedback, breaking the ranking and updating pairs independently can be more computationally efficient—and even more stable—than learning directly from the PL loss. Moreover, this provides a theoretical explanation for the empirical success of RLHF in LLMs (e.g., ouyang2022training), where ranking feedback is decomposed into pairwise comparisons for parameter estimation.
@@ 5.2 Lower Bound
Suppose $T\geq d^{2}/(8K^{2})$. Define the feature space as $\Phi:=\mathcal{S}^{d-1}$, the unit sphere in $\mathbb{R}^{d}$, and let the parameter space be $\Theta=\{-\mu,\mu\}^{d}$, where $\mu=\sqrt{d/(8K^{2}T)}$. Then, for any final policy $\widehat{\pi}_{T}\in\triangle_{\Phi}$ returned after collecting $T$ samples (using any sampling policy), the expected suboptimality gap is lower bounded as:
@@
Discussion of Theorem 3. The proof is deferred to Appendix F. Theorem 3 provides theoretical support for our upper bounds, especially in their dependence on $K$; in particular, it matches the upper bound in Theorem 2 when $|S_{t}|=K$. To the best of our knowledge, this is the first lower bound on the suboptimality gap that incorporates PL ranking feedback in PbRL and formally shows that the suboptimality gap can diminish as $K$ grows, highlighting the advantage of utilizing ranking feedback over simple pairwise comparisons.
@@ 6 Numerical Experiments
We conduct two sets of experiments to empirically validate our theoretical findings: (i) one using synthetic data (Subsection 6.1), and (ii) another using two real-world datasets (Subsection 6.2). We compare our proposed algorithm, M-AUPO, against three baselines: (i) DopeWolfe [thekumparampil2024comparing], which selects $K$ actions in a non-contextual setting; (ii) Uniform, which uniformly samples assortments of size $K$ at random; and (iii) Best&Ref constructs an action pair ($|S_{t}|=2$) by combining the action that maximizes the current reward estimate with another sampled from a reference policy (e.g., uniform random or SFT), following the setup in Online GSHF [xiong2023iterative] and XPO [xie2024exploratory]. In our experiments, the reference policy for Best&Ref is set to the uniform random policy.
@@ 6.1 Synthetic Data
The first two plots in Figure 1 show the suboptimality gap of M-AUPO under both the PL loss (3) and RB loss (4) as the maximum assortment size $K$ varies. The results clearly show that performance improves as $K$ increases, supporting our theoretical findings. In the third plot of Figure 1, we compare the performance of M-AUPO with three baseline methods at the final round for $K=5$, showing that our algorithm significantly outperforms all baselines. While DopeWolfe also considers the selection of $K$ actions from $N$ actions, it treats each context $x$ independently and is specifically designed for the context-free setting (i.e., a singleton context). As a result, DopeWolfe cannot leverage information sharing across varying contexts and performs poorly in our setting. Furthermore, M-AUPO outperforms naive assortment selection strategies such as Uniform and Best&Ref, as it explicitly chooses assortments that maximize the average uncertainty, thereby achieving more efficient exploration. Finally, the results also show that the RB version of M-AUPO empirically outperforms the PL-based approach, since RB does not incur approximation errors when computing expectations. See Appendix H.1 for additional experimental results.
@@ 6.2 Real-World Dataset
We use the gemma-2b [^7] https://huggingface.co/google/gemma-2b-it [team2024gemma] LLM to construct the feature $\phi(x,a)$. Specifically, $\phi(x,a)$ is obtained by extracting the embedding of the concatenated prompt and response from the last hidden layer of the LLM, with size $d=2048$. Additionally, we use the Mistral-7B [jiang2023mistral7b] reward model [^8] https://huggingface.co/Ray2333/reward-model-Mistral-7B-instruct-Unified-Feedback as the true reward model $r_{\bm{\theta}^{\star}}$ to generate ranking feedback and compute the suboptimality gap accordingly. We measure the suboptimality gap every 2,500 rounds and report the average over 10 independent runs, along with the standard error. In these experiments, we report results only for the RB version of M-AUPO, as it is computationally more efficient and empirically performs better, as shown in Figure 1.
The first two plots in Figure 2 show that the performance improves as $K$ increases on two real-world datasets In the third plot of Figure 2, we compare the performance of M-AUPO with other baselines with $K=3$ at the final round, showing that M-AUPO outperforms baselines by a large margin. See Appendix H.2 for additional experimental details and results.
Figure 1: Synthetic data experiment: suboptimality gap of M-AUPO under varying $K$, evaluated with PL loss (left) and RB loss (middle), along with comparison against other baselines (right).
Figure 2: Real-world dataset experiment: suboptimality gap of M-AUPO under varying $K$ on the TREC-DL dataset (left) and the NECTAR dataset (middle), along with comparison against other baselines (right). The results are rescaled to align the performances between the two datasets.
@@ 7 Conclusion
To the best of our knowledge, this work provides the first theoretical result in online PbRL showing that the suboptimality gap decreases as more options are revealed to the labeler for ranking feedback. Our analysis also removes the $\mathcal{O}(e^{B})$ dependency in the leading term without modifying the algorithm, implying that existing PbRL and dueling bandit methods can similarly avoid this dependence through our refined analysis. Together, these results advance the theoretical understanding of PbRL, revealing both the value of richer feedback and the opportunity for more refined and efficient analysis.
@@ Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2022-NR071853, RS-2023-00222663, and RS-2025-25420849), by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2025-02263754), by the AI-Bio Research Grant through Seoul National University, and by the 2025 Global Google PhD Fellowship funded with support from Google.org.
Part Appendix
@@ Appendix A Further Related Work
Logistic and MNL bandits. Our work is also closely related to logistic bandits and multinomial logit (MNL) bandits. The logistic bandit problem [dong2019performance, faury2020improved, abeille2021instance, faury2022jointly, lee2024improved, lee2024unified] is a special case of the MNL bandit model in which the agent offers only a single item (i.e., $K=1$) at each round and receives binary feedback indicating whether the item was selected (1) or not (0). faury2020improved examined how the regret in logistic bandits depends on the non-linearity parameter $\kappa$ of the logistic link function and proposed the first algorithm whose regret bound eliminates explicit dependence on $1/\kappa=\mathcal{O}(e^{B})$. abeille2021instance further improved the theoretical dependency on $1/\kappa$ and established a matching, problem-dependent lower bound. Building on this, faury2022jointly developed a computationally efficient algorithm whose regret still matches the lower bound established by abeille2021instance.
@@
Multinomial logit (MNL) bandits tackle a more sophisticated problem than logistic bandits. Instead of offering a single item and observing binary feedback, the learner chooses a subset of items—underscoring the combinatorial nature of the task—and receives non-uniform rewards driven by an MNL choice model [agrawal2019mnl, agrawal2017thompson, ou2018multinomial, chen2020dynamic, oh2019thompson, oh2021multinomial, perivier2022dynamic, agrawal2023tractable, lee2024nearly, lee2025improved]. A recent breakthrough by lee2024nearly closed a long-standing gap by providing a computationally efficient algorithm that attains the minimax-optimal regret for this setting. Building on this result, lee2025improved further reduced the regret bound by a factor polynomial in $B$ and logarithmic in $K$, and established the first variance-dependent regret bounds for MNL bandits.
@@
Our work extends the online confidence bound analysis of lee2025improved to the Plackett–Luce (PL) model. This extension is natural because the PL probability distribution decomposes into a sequence of MNL probabilities over successive choices. Crucially, we leverage their key insight—that the MNL loss exhibits an $\ell_{\infty}$-self-concordant property—to eliminate the harmful $\mathcal{O}(e^{B})$ dependence. This is one of the main contributions of our work (see Lemma C.2).
@@
RL with MNL models. Recent work has extended the Multinomial Logit (MNL) framework beyond bandit formulations to reinforcement learning. lee2025combinatorial introduced combinatorial RL with preference feedback, a framework in which an agent learns to select subsets of items so as to maximize long-term cumulative rewards.
@@
Another line of research incorporates MNL models directly into the transition dynamics. hwang2022model proposed MNL-MDPs, a class of Markov decision processes whose transition probabilities follow an MNL parameterization. Building upon this formulation, cho2024randomized improved the regret bounds by improving the exponential dependence on $B$, and park2025infinite extended the analysis to the infinite-horizon setting.
Appendix B Notation
@@
Let $T$ denote the total number of rounds, with $t\in[T]$ representing the current round. We use $N$ for the total number of items, $K$ for the maximum assortment size, $d$ for the feature vector dimension, and $B$ as an upper bound on the norm of the unknown parameter. For notational convenience, we provide Table B.1.
@@
${\mathcal{T}}^{w}$ $\left\{t\in[T]:\max_{a,a^{\prime}\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|_{H_{t}^{-1}}\geq\frac{1}{3\sqrt{2}K\beta_{T+1}(\delta)}\right\}$, warm-up rounds
$M_{t}$ $\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{s}|}\mathbb{E}_{\sigma\sim\mathbb{P}_{s}(\cdot|S_{s},\bm{\theta}^{\star})}\left[\nabla^{2}\ell^{(j)}_{s}(\widehat{\bm{\theta}}_{s})\right]+\lambda\mathbf{I}_{d}$
@@
Appendix C Proof of Theorem 1